The national park is in a circular shape. The middle portion has playthings for kids surrounded by a circular walking pathway, then calculate the walking area whose outer radius is 18m

 and the inner radius is 7m.



1
Expert's answer
2021-01-12T01:06:37-0500

The walking area whose outer radius is 18m and the inner radius is 7m is:

"A = \pi×R1^2 - \pi×R2^2 = \pi×(18^2 - 7^2) = 275\pi = 863.94 m^2."